A Yoplait factory produced 375,000 cups of yogurt yesterday. The factory has a rated capacity of 410,000 cups per day and the plant manager estimates that the most the factory could produce in one day (its theoretical capacity) is 520,000 cups of yogurt. Product is loaded on trucks for customers and the standard time to load a truck is 120 minutes.

What was the efficiency of the factory yesterday?

What was the utilization of the factory yesterday?

If a truck is loaded in 100 minutes, what was the efficiency of the loading operation?
Answer #1

1) According to rated capacity, efficiency can be calculated as

375000 / 410000 = 0.9146
or 91.46%

Operating efficiency is 91.46%

2) The factory has produced 375000 units while its theoretical capacity is 520000

375000 / 520000 = 0.7211
or 72.11%

Capacity utilization is 72.11%

3) Standard time to load the truck is 120 minutes while it has been loaded in 100 minutes actually.

100 / 120 = 0.8333
or 83.33%

Truck loading efficiency is 83.33%